OpenStudy (anonymous):

Find the counterexample to this statement: If the product of two numbers is even, then the two numbers are even. Answer (A) 2·4 = 8 (B) 3·5 = 15 (C) 30·2·2 = 120 (D) 3·4 = 12

OpenStudy (anonymous):

Well you want the one where the product is even but one of the numbers is odd. An example that's not one of the answers would be 3*6=18

OpenStudy (anonymous):

Remember a conditional statement is only false when the If clause is true and the Then clause is false

OpenStudy (zzr0ck3r):

P->~Q

OpenStudy (anonymous):

IDK it has to be one of the answers Thanks for the Help

OpenStudy (btaylor):

D is the right answer, because 3 is not even.

mathslover (mathslover):

by the way @angelopps : \[\huge{\mathbb{Welcome}\textbf{To}\mathbb{Open}\textbf{Study}}\]

OpenStudy (anonymous):

Thanks @btaylor and everyone else